# Accurate Pocket Identification for Binding-Site-Agnostic Docking

### v1

@@ Abstract
Accurate identification of druggable pockets is essential for structure-based drug design. However, most pocket-identification algorithms prioritize their geometric properties over downstream docking performance. To address this limitation, we developed RAPID-Net, a pocket-finding algorithm for seamless integration with docking workflows. When guiding AutoDock Vina, RAPID-Net outperforms DiffBindFR on the PoseBusters benchmark and enables blind docking on large proteins that AlphaFold 3 cannot process as a whole. Furthermore, RAPID-Net surpasses PUResNet and Kalasanty in docking accuracy and pocket-ligand intersection rates across diverse datasets, including PoseBusters, Astex Diverse Set, BU48, and Coach420. When accuracy is evaluated as “at least one correct pose in the ensemble”, RAPID-Net outperforms AlphaFold 3 on the PoseBusters benchmark, suggesting that our approach can be further improved with a suitable pose reweighting tool offering a cost-effective and competitive alternative to AlphaFold 3 for docking. Finally, using several therapeutically relevant examples, we demonstrate the ability of RAPID-Net to identify remote functional sites, highlighting its potential to facilitate the development of innovative therapeutics.
Our model is publicly available: https://github.com/BalytskyiJaroslaw/RAPID-Net
@@ I Introduction and Problem Statement
Molecular docking is the computational problem of predicting the most favorable ligand binding poses in a protein-ligand complex, given the experimentally determined or computer-simulated protein structure and the initial conformation of the ligand 1. It plays a key role in drug development, typically following target identification when the biological molecule responsible for a disease is identified and validated as a therapeutic target 2. Once the target is identified, molecular docking is used for structure-based drug discovery and development 3; 4 to identify and optimize potential lead compounds 5; 6. The docking results can guide virtual screening workflows, enabling the selection of promising candidates from vast chemical libraries 7; 8. This process helps to determine whether a compound has potential for drug development 9. By significantly improving the speed and efficiency of early drug discovery, molecular docking reduces the time and costs associated with the development of new therapeutics 10.
Typical protein-ligand docking pipelines rely on users specifying the binding pocket, and docking programs such as AutoDock Vina 11; 12, GOLD 13, and Glide 14, use grids to confine their search to known or hypothesized protein’s interaction sites. However, in the absence of such information – in “blind” or binding-site-agnostic settings – protein-ligand docking becomes significantly more difficult 15, as the docking algorithm must scan the entire protein surface, dramatically increasing the computational complexity. Traditional blind docking methods often rely on extensive sampling to explore potential binding sites across the whole protein, but this approach is computationally and time-intensive, making it impractical for large-scale virtual screening tasks 16; 17.
@@
With the advent of advanced protein structure prediction methods 18–including AlphaFold 19, ColabFold 20, OpenFold 21, and RosettaFold 22–the number of generated protein structures has surged, often without any corresponding ligand information. Meanwhile, only about $\sim 5\%$ of human coding genes currently serve as commercial drug targets, leaving a vast array of disease-related targets unexplored 23. Despite this immense potential, most new drug research and development remains focused on a limited set of well-established targets, underscoring the urgent need for novel druggable targets 24. As more proteins lacking known binding pocket information are considered in drug discovery 25, the demand for ligand docking approaches that do not rely on prior knowledge of the binding site has grown significantly, highlighting the necessity for binding-site-agnostic methods.
At the same time, many existing protein binding site identification algorithms are evaluated primarily on the geometric properties of their predicted pockets 26, rather than on how accurately they guide subsequent ligand docking. In this work, our primary objective is to develop a pocket-finding algorithm that can be readily integrated with standard docking software and provide accurate, efficient guidance in binding-site-agnostic settings. We evaluate its performance both in terms of docking accuracy–using AutoDock Vina 1.2.5 12–and the geometric characteristics of the generated pockets.
@@
In terms of underlying pocket prediction function, existing algorithms can be broadly classified into classical and Machine Learning (ML)-driven approaches 26. Classical methods rely on expert-defined rules or heuristics to detect pockets, whereas ML-driven algorithms learn to extract features from protein data without explicit human instructions. In this work, we adopt the definition of Machine Learning by François Chollet as “the effort to automate intellectual tasks normally performed by humans” 27. From this perspective, the main objective of an ML framework is to unveil a meaningful data representation that allows pocket-prediction rules to emerge automatically, rather than being manually hardcoded.
@@
Furthermore, “Deep Learning (DL) is a specific subfield of Machine Learning: a new take on learning representations from data that puts an emphasis on learning successive layers of increasingly meaningful representations” 27. In the context of pocket identification in our study, these layers of data representations are realized through Deep Neural Networks (DNNs). Finally, in contrast to DL, so-called “shallow” learning approaches typically employ only one or a small number of consecutive data representation layers 27.
@@
In the following Sections, we describe our DL model for pocket identification, named RAPID-Net (ReLU Activated Pocket Identification for Docking). We tested RAPID-Net on various benchmarks to demonstrate its efficiency and analyzed its predictions on several therapeutically important proteins. In Section II, we discuss the architecture and design rationale behind our model. In Section III, we present the RAPID-Net training pipeline, highlighting its key differences from existing approaches, including the use of ReLU activation in the last layer of the network to operate on a “soft” rather than binary labels. Using the RAPID-Net model developed in Sections II and III, in Section IV we integrate it into a docking protocol that is subsequently used for docking. In Section V, we describe the evaluation metrics used to assess the quality of the model’s predictions. Sections VI and VII report docking results on the PoseBusters 28 and Astex Diverse Set 29 datasets, respectively. For completeness and to provide a direct side-by-side comparison with existing pocket prediction algorithms, Section VIII evaluates the geometric characteristics of predicted pockets on the Coach420 30 and BU48 31 datasets. To demonstrate the ability of RAPID-Net to identify distant binding sites, Section IX analyzes its predictions for several therapeutically important proteins with known distal sites. Finally, in Section X, we reach our conclusions and outline the potential directions for future work.
II Design Rationale behind our Model
Figure 1: Alpha-spheres predicted by FPocket 32 and the majority-voted pocket predicted by our RAPID-Net model for the ABHD5 protein.
@@ II Design Rationale behind our Model
Figure 3: For the 8F4J protein structure from the PoseBusters 28 dataset, our docking approach passes the RMSD test and 25/26 PoseBusters 28 validation checks, outperforming AlphaFold 3 33, which cannot process the structure as a whole. Unlike residue-focused methods such as PUResNet V2 34, which produce complex predictions that are difficult to interpret, our approach offers clear, plug-and-play guidance for docking grid selection, as illustrated in the bottom panel.
@@
When predicting likely binding pockets, there is a well-known trade-off between precision and recall. A recent evaluation of pocket prediction methods based on the geometrical properties of their predicted pockets 26 highlights this trade-off. Currently available ML methods such as VN-EGNN 35, GrASP 36, and PUResNet 37 achieve high precision (over 90%), but they systematically predict a small number of pockets, leading to low recall. As 26 points out, generating multiple predictions, some of which may be false positives, is often more useful than potentially missing viable binding sites.
@@
As demonstrated on diverse benchmark datasets in the following Sections, including PoseBusters 28, Astex Diverse Set 29, Coach420 30, and BU48 31, the proposed ensemble-based approach yields more stable and reliable performance compared to current ML-driven pocket predictors, both considering the majority-voted and minority-reported pockets.
@@
On the other hand, FPocket 32 typically predicts many, often overlapping alpha-spheres, making it difficult or even impossible to segment the protein surface clearly. FPocket 32 can place alpha spheres in adjacent but distinct pockets, introducing ambiguity in the definition of distinct docking regions. As an illustration, for the ABHD5 protein structure obtained from AlphaFold 19 shown in Fig. 1, the majority-voted pocket predicted by our model forms a compact clearly defined region. If FPocket 32 was used in the docking pipeline instead, nearly the entire protein surface would need to be included, drastically increasing computational costs.
The drawbacks of overly large search grids can be illustrated by the 8F4J protein structure from the PoseBusters 28 (PB) dataset, shown in the top panel of Fig. 3. Even advanced docking platforms such as AlphaFold 3 33, which incorporate FPocket 32 to locate potential docking sites, cannot process the entire 8F4J protein as a whole: “Another PDB entry (8F4J) was too large to inference the entire system (over 5,120 tokens), so we included only protein chains within $20\ \text{\AA}$ of the ligand of interest.” 33. In contrast, as shown in Fig. 3 and discussed further in the text, the RAPID-Net guidance allows AutoDock Vina 11; 12to generate a pose passing the $\text{RMSD}<2\ \text{\AA}$ test and satisfy 25/26 PoseBusters (PB) tests.
@@
Kalasanty 38 and PUResNet V1111PUResNet V137 predicts cavities where the ligand is likely to reside, whereas PUResNet V2 34 predicts residues likely to interact with the ligand. For simplicity, we refer to PUResNet V1 as “PUResNet” throughout the text, as in this work we focus on cavities for generating an accurate search grid for subsequent docking. 37 both detect potential ligand-binding pockets using a voxel-based representation, where each voxel is $2\text{\AA}\times 2\text{\AA}\times 2\text{\AA}$ in size.
@@
In contrast, PRANK 39 and DeepPocket 40 reweight the alpha-spheres identified by the rule-based FPocket 32 algorithm. As a result, these methods inherit FPocket’s initial search space and any associated cascading errors. Moreover, alpha-sphere-based approaches may lack sufficient granularity and can struggle to capture subtle variations among binding sites.
@@
GrASP 36, IF-SitePred 41, VN-EGNN 35, and PUResNetV2 34 provide predictions of potential ligand interactions at the residue level. Although these predictions can reveal important binding details, they are less suitable for docking workflows that require a well-defined three-dimensional region for accurate ligand placement and orientation. By precisely defining these regions, the computational search space is reduced as docking is restricted to reasonable binding pockets. As illustrated in the bottom panel of Fig. 3, residue-level predictions can be difficult to interpret when defining a reasonable search grid.
@@
Although some studies reweight residues to guide blind docking 42, and others 25 employ voxelized pocket prediction models–such as PUResNet 37–in tandem with AutoDock Vina 12, our work focuses on developing a voxel-based cavity prediction model that seamlessly integrates into the docking pipeline. This voxel-based approach not only provides well-defined search regions but also facilitates a more modular and interpretable workflow compared to less structured docking strategies, as illustrated in Fig. 5 and discussed further in Section IV.
@@
Furthermore, as discussed in Section VI, improving the accuracy of pocket identification directly improves docking results. In particular, AutoDock Vina 12, guided by our pocket predictor, outperforms the state-of-the-art DiffBindFR 43 docking tool, which otherwise scans the entire protein surface in “blind” settings.
@@
Our model architecture. Fig. 4 depicts the architecture of our proposed RAPID-Net model, which is similar to U-Net 44 with encoder and decoder branches. However, we implement several notable adaptations.
@@
First, we include residual connections only in the encoder part of the model, as our experiments show that they are highly beneficial there, but lead to overfitting if included in the decoder. This approach differs from Kalasanty 38, which omits residual blocks altogether 45, and PUResNet 37, which uses them throughout the network.
@@
Second, although the standard SE-ResNet 46 architecture typically includes attention blocks at multiple layers, studies on breast-cancer imaging 47 and Raman spectra classification 48 have shown that using a single attention block can be highly effective while mitigating overfitting. Our experiments similarly suggest that in inherently noisy datasets, adding too many attention modules can amplify noise and degrade performance.
@@
III Model Training and Inference Pipeline
We trained our model using the sc-PDB 49 dataset, which contains protein structures with annotated binding sites. In this dataset, cavities are defined using VolSite 50, which maps the pharmacophore properties of nearby protein atoms onto a three-dimensional grid. This method assumes that ligand-induced conformational changes remain relatively small 51, treating each cavity atom as a “pseudoatom” to denote an interaction point rather than a physical atom.
Following 37, we used a curated subset of sc-PDB 49 in which redundant protein structures were filtered out based on their Tanimoto index 52. Each protein structure was then placed in a $36\times 36\times 36$ grid, with each voxel corresponding to a $2\text{\AA}\times 2\text{\AA}\times 2\text{\AA}$ unit cell. We extracted 18 features per voxel using the tfbio 53 package to provide a representation of the protein environment for the model training.
In contrast to previous studies, we apply the medical image segmentation practices 54; 55; 56 to perform soft segmentation of ligand-binding pockets, rather than treating it as a strict binary classification problem. In Kalasanty 38 and PUResNet 37, each voxel is considered part of a binding pocket if it contains at least one cavity pseudoatom. The number of pseudoatoms per voxel is then clipped to the range $[0,1]$, and the model is trained with a Dice loss function as part of a binary classification task:

### v2

# RAPID-Net: Accurate Pocket Identification for Binding-Site-Agnostic Docking

@@ Abstract
Accurate identification of druggable pockets and their features is essential for structure-based drug design and effective downstream docking. Here, we present RAPID-Net, a deep learning-based algorithm designed for the accurate prediction of binding pockets and seamless integration with docking pipelines. On the PoseBusters benchmark, RAPID-Net–guided AutoDock Vina achieves 54.9% of Top-1 poses with RMSD $<2\,\text{\AA}$ and satisfying the PoseBusters chemical‐validity criterion, compared to 49.1% for DiffBindFR. On the most challenging time split of PoseBusters aiming to assess generalization ability (structures submitted after September 30, 2021), RAPID-Net-guided AutoDock Vina achieves 53.1% of Top-1 poses with RMSD $<2\,\text{\AA}$ and PB-valid, versus 59.5% for AlphaFold 3. Notably, in 92.2% of cases, RAPID-Net-guided Vina samples at least one pose with RMSD $<2\,\text{\AA}$ (regardless of its rank), indicating that pose ranking, rather than sampling, is the primary accuracy bottleneck. The lightweight inference, scalability, and competitive accuracy of RAPID-Net position it as a viable option for large-scale virtual screening campaigns. Across diverse benchmark datasets, RAPID-Net outperforms other pocket prediction tools, including PUResNet and Kalasanty, in both docking accuracy and pocket–ligand intersection rates. Furthermore, we demonstrate the potential of RAPID-Net to accelerate the development of novel therapeutics by highlighting its performance on pharmacologically relevant targets. RAPID-Net accurately identifies distal functional sites, offering new opportunities for allosteric inhibitor design. In the case of the RNA-dependent RNA polymerase of SARS-CoV-2, RAPID-Net uncovers a wider array of potential binding pockets than existing predictors, which typically annotate only the orthosteric pocket and overlook secondary cavities.
RAPID-Net, along with interactive demonstration notebooks, is publicly available at: github.com/BalytskyiJaroslaw/RAPID-Net, offering a user-friendly and fully reproducible framework for pocket-guided docking.
@@ I Introduction and Problem Statement
Molecular docking is the computational problem of predicting the most favorable ligand binding poses in a protein-ligand complex, given the experimentally determined or computer-simulated protein structure and the initial conformation of the ligand 1. It plays a key role in drug development, typically following target identification when the biological molecule responsible for a disease is identified and validated for therapeutic intervention 2. Once the target is identified, molecular docking is used for structure-based drug discovery and development 3; 4 to identify and optimize potential lead compounds 5; 6. The docking results can guide virtual screening workflows, enabling the selection of promising candidates from vast chemical libraries 7; 8. This process helps to determine whether a compound has potential for drug development 9. By significantly improving the speed and efficiency of early drug discovery, molecular docking reduces the time and costs associated with the development of new therapeutics 10.
@@
Typical protein-ligand docking pipelines rely on users specifying the binding pocket, and docking programs such as AutoDock Vina 11; 12, GOLD 13, and Glide 14 use grids to confine their search to known or hypothesized protein’s interaction sites. However, in the absence of such information – in “blind” or binding-site-agnostic settings – protein-ligand docking becomes significantly more difficult 15, as the docking algorithm must scan the entire protein surface, dramatically increasing the computational complexity. Traditional blind docking methods often rely on extensive sampling to explore potential binding sites across the whole protein, but this approach is computationally and time-intensive, making it impractical for large-scale virtual screening tasks 16; 17. With the advent of advanced protein structure prediction methods 18–including AlphaFold 19, ColabFold 20, OpenFold 21, and RosettaFold 22–the number of protein structures generated has surged, often without any corresponding ligand information. Meanwhile, only about $\sim 5\%$ of the human coding genes currently serve as commercial drug targets, leaving a wide range of disease-related targets unexplored 23. Despite this immense potential, most new drug research and development remain focused on a limited set of well-established targets, underscoring the urgent need for novel druggable targets 24. As more proteins lacking known binding pocket information are considered in drug discovery 25, the demand for ligand docking approaches that do not rely on prior knowledge of the binding site has grown significantly, highlighting the necessity for binding-site-agnostic methods.
In this work, we aim to develop a high‐accuracy pocket-finding algorithm (RAPID-Net) optimized for seamless integration with standard docking pipelines to deliver efficient guidance in binding-site-agnostic workflows. Many protein binding site predictors are benchmarked solely on the geometric properties of their predicted pockets 26, but their impact on the accuracy of downstream ligand docking is rarely assessed. While many factors affect docking performance, we show further in the text that pocket prediction quality has a decisive role in determining docking success. We evaluate RAPID-Net’s performance both by docking accuracy – using the widely adopted AutoDock Vina 12 – and by pocket-ligand intersection rates, as detailed further in the text. Indeed, RAPID-Net can provide the search grid centers and dimensions for any docking engine, and Vina was chosen due to its popularity.
Despite relying on Vina’s relatively simple scoring function, RAPID-Net-guided docking outperforms the state-of-the-art blind-docking tool DiffBindFR 27 by a substantial margin (54.9% vs 49.1%) on the PoseBusters benchmark 28 (see Section VII). This result underscores our central claim: precise pocket identification is a decisive driver of downstream docking success. It is important to note that, unlike DiffBindFR, which incorporates the ligand-induced protein changes, in our approach, we use our pocket predictor to guide AutoDock Vina that docks to a rigid receptor, and all evaluations are performed on holo structures. Nevertheless, the same pocket and search grid localization can boost flexible-receptor engines as well, for example, DSDPFlex 29:“By leveraging prior knowledge or information, this innovative approach is anticipated to enhance the search process within the appropriate binding pocket.”. Overall, our results indicate that accurate pocket localization currently outweighs receptor flexibility as the key bottleneck, making RAPID-Net a practical route to higher accuracy in binding-site-agnostic workflows and a solid foundation on which ligand-induced conformational refinements can later be added as a subleading effect. We refer the treatment of ligand-induced conformational changes to our future work, focusing here on maximizing the pocket prediction accuracy.
For comparison with co-folding approaches such as AlphaFold 3 (AF3) 30, we follow the PoseBench protocol 31. This protocol uses a curated subset of 130 protein-ligand complexes from PoseBusters 28 deposited after September 30, 2021, specifically designed to evaluate generalization performance rather than memorization 31. On this challenging benchmark, RAPID‑Net + Vina achieves a Top-1 success rate of 53.1%, only six percentage points below AF3’s 59.5% (see Section VII), while requiring substantially fewer computational resources and offering significantly better scalability for virtual screening.
RAPID-Net supplies an accurate search box for subsequent lightweight rigid-receptor redocking, whereas AF3 predicts the full protein-ligand complex from the protein sequence and the ligand SMILES. The significantly lower computational demand of our approach is exemplified by the 8F4J structure 28, which AF3 was unable to process as a whole: “Another PDB entry (8F4J) was too large to inference the entire system (over 5,120 tokens), so we included only protein chains within $20\ \rm{\AA}$ of the ligand of interest.”. In contrast, RAPID-Net guided Vina to a pose with $\mathrm{RMSD}<2\,\text{\AA}$ [^1] A demonstration video is available at: https://youtu.be/EkUKmoW11pE?si=3aKBkL3ZRq8ibWqo, as illustrated in Fig. 2. Thus, RAPID-Net provides near-state-of-the-art accuracy while remaining practical for large-scale virtual screening.
@@
In terms of the underlying pocket prediction function, existing pocket predicting algorithms can be broadly classified into classical and Machine Learning (ML)-driven approaches 26. Classical methods rely on expert-defined rules or heuristics to detect pockets, whereas ML-driven algorithms learn to extract features from protein data without explicit human instructions. In this work, we adopt the definition of Machine Learning by François Chollet as “the effort to automate intellectual tasks normally performed by humans” 32. From this perspective, the main objective of an ML framework is to unveil a meaningful data representation that allows pocket-prediction rules to emerge automatically, rather than being manually hardcoded.
@@
Furthermore, “Deep Learning (DL) is a specific subfield of Machine Learning: a new take on learning representations from data that puts an emphasis on learning successive layers of increasingly meaningful representations” 32. In the context of pocket identification in our study, these layers of data representations are realized through Deep Neural Networks (DNNs). Finally, in contrast to DL, so-called “shallow” learning approaches typically employ only one or a small number of consecutive data representation layers 32.
@@
In the following Sections, we describe our DL model for pocket identification, named RAPID-Net (ReLU Activated Pocket Identification for Docking). We tested RAPID-Net on various benchmarks to demonstrate its efficiency and analyzed its predictions on several therapeutically important proteins. In Section II, we provide a brief overview of existing pocket prediction algorithms. In Section III, we discuss the architecture and design rationale behind our model. In Section IV, we present the RAPID-Net training pipeline, highlighting its key differences from existing approaches, including the use of ReLU activation in the last layer of the network to operate on a “soft” rather than binary labels. Using the RAPID-Net model developed in Sections III and IV, in Section V we integrate it into a docking protocol that is subsequently used for docking. In Section VI, we describe the evaluation metrics used to assess the quality of the model’s predictions. Sections VII and VIII report docking results on the PoseBusters 28 and Astex Diverse Set 33 datasets, respectively. For completeness and to provide a direct side-by-side comparison with existing pocket prediction algorithms, Section IX evaluates the geometric characteristics of predicted pockets on the Coach420 34 and BU48 35 datasets. To demonstrate the ability of RAPID-Net to identify distant binding sites, Section X analyzes its predictions for several therapeutically important proteins with known distal sites. Finally, in Section XI, we reach our conclusions and outline the potential directions for future work.
II Overview of the Available Pocket Prediction Algorithms
Pocket prediction methods exploit a variety of different techniques to identify potential binding sites. Geometry-based techniques like Fpocket 36, Ligsite 37 and Surfnet 38 identify cavities by analyzing the geometry of the molecular surface of a protein and usually rely on the use of a grid, gaps, spheres, or tessellation 36; 37; 38; 39; 40; 41; 42; 43. Energy-based methods such as PocketFinder 44 rely on the calculation of interaction energies between the protein and chemical group or probe to identify cavities 44; 45; 46; 47; 48; 49 .
Conservation-based methods make use of sequence evolutionary conservation information to find patterns in multiple sequence alignments and identify conserved key residues for ligand binding site identification 50; 51; 52.
Template-based methods rely on structural information from homologues and the assumption that structurally conserved proteins might bind ligands at a similar location 53; 54; 34; 55; 56; 57. Combined approaches or meta-predictors combine multiple methods, or the use of multiple types of data to infer ligand binding sites, e.g., geometric features with sequence conservation 35; 58; 59; 60; 61; 62; 63.
Although classical geometry-based tools such as Fpocket are fast and widely applicable, they can sometimes merge individual pockets into a single “blurred” cavity, for example, as observed in key SARS-CoV-2 targets 64; 65. In Section X, we revisit the Nsp12 protein considered in 64; 65 using the RAPID-Net model we developed to demonstrate its ability to identify functionally significant sites.
III Design Rationale behind our Model
@@
Figure 2: For the 8F4J protein structure from the PoseBusters 28 dataset, our docking approach achieves $\text{RMSD}<2$, while AlphaFold 3 30 is unable to process the full structure due to input size limitations. Unlike residue-focused methods such as PUResNet V2 66, which produce complex predictions that are difficult to interpret, our approach offers clear, plug-and-play guidance for docking grid selection, as illustrated in the bottom panel.
@@
When predicting likely binding pockets, there is a well-known trade-off between precision and recall. A recent evaluation of pocket prediction methods based on the geometrical properties of their predicted pockets 26 highlights this trade-off. Currently available ML methods such as VN-EGNN 67, GrASP 68, and PUResNet 69 achieve high precision (over 90%), but they systematically predict a small number of pockets, leading to low recall. As 26 points out, generating multiple predictions, some of which may be false positives, is often more useful than potentially missing viable binding sites.
@@
As demonstrated on diverse benchmark datasets in the following Sections, including PoseBusters 28, Astex Diverse Set 33, Coach420 34, and BU48 35, the proposed ensemble-based approach yields more stable and reliable performance compared to current ML-driven pocket predictors, both considering the majority-voted and minority-reported pockets.
@@
Kalasanty 70 and PUResNet V1222PUResNet V169 predicts cavities where the ligand is likely to reside, whereas PUResNet V2 66 predicts residues likely to interact with the ligand. For simplicity, we refer to PUResNet V1 as “PUResNet” throughout the text, as in this work we focus on cavities for generating an accurate search grid for subsequent docking. 69 both detect potential ligand-binding pockets using a voxel-based representation, where each voxel is $2\text{\AA}\times 2\text{\AA}\times 2\text{\AA}$ in size.
@@
In contrast, PRANK 71 and DeepPocket 72 reweight the alpha-spheres identified by the rule-based FPocket 73 algorithm. As a result, these methods inherit FPocket’s initial search space and any associated cascading errors. Moreover, alpha-sphere-based approaches may lack sufficient granularity and can struggle to capture subtle variations among binding sites, as previously noted 64; 65.
@@
GrASP 68, IF-SitePred 74, VN-EGNN 67, and PUResNetV2 66 provide predictions of potential ligand interactions at the residue level. Although these predictions can reveal important binding details, they are less suitable for docking workflows that require a well-defined three-dimensional region for accurate ligand placement and orientation. By precisely defining these regions, the computational search space is reduced as docking is restricted to reasonable binding pockets. As illustrated in the bottom panel of Fig. 2, residue-level predictions can be difficult to interpret when defining a reasonable search grid.
@@
Although some studies reweight residues to guide blind docking 75, and others 25 employ voxelized pocket prediction models–such as PUResNet 69–in tandem with AutoDock Vina 12, our work focuses on developing a voxel-based cavity prediction model that seamlessly integrates into the docking pipeline. This voxel-based approach not only provides well-defined search regions but also facilitates a more modular and interpretable workflow compared to less structured docking strategies, as illustrated in Fig. 4 and discussed further in Section V.
@@
Furthermore, as discussed in Section VII, improving the accuracy of pocket identification directly improves docking results. In particular, AutoDock Vina 12, guided by our pocket predictor, outperforms the state-of-the-art DiffBindFR 27 docking tool, which otherwise scans the entire protein surface in “blind” settings.
@@
Our model architecture. Fig. 3 depicts the architecture of our proposed RAPID-Net model, which is similar to U-Net 76 with encoder and decoder branches. However, we implement several notable adaptations.
@@
First, we include residual connections only in the encoder part of the model, as our experiments show that they are highly beneficial there, but lead to overfitting if included in the decoder. This approach differs from Kalasanty 70, which omits residual blocks altogether 77, and PUResNet 69, which uses them throughout the network.
@@
Second, although the standard SE-ResNet 78 architecture typically includes attention blocks at multiple layers, studies on breast-cancer imaging 79 and Raman spectra classification 80 have shown that using a single attention block can be highly effective while mitigating overfitting. Our experiments similarly suggest that in inherently noisy datasets, adding too many attention modules can amplify noise and degrade performance.
@@
IV Model Training and Inference Pipeline
We trained our model using the sc-PDB 81 dataset, which contains protein structures with annotated binding sites. In this dataset, cavities are defined using VolSite 82, which maps the pharmacophore properties of nearby protein atoms onto a three-dimensional grid. This method assumes that ligand-induced conformational changes remain relatively small 83, treating each cavity atom as a “pseudoatom” to denote an interaction point rather than a physical atom.
Following 69, we used a curated subset of sc-PDB 81 in which redundant protein structures were filtered out based on their Tanimoto index 84. Each protein structure was then placed in a $36\times 36\times 36$ grid, with each voxel corresponding to a $2\text{\AA}\times 2\text{\AA}\times 2\text{\AA}$ unit cell. We extracted 18 features per voxel using the tfbio 85 package to provide a representation of the protein environment for the model training.
In contrast to previous studies, we apply the medical image segmentation practices 86; 87; 88 to perform soft segmentation of ligand-binding pockets, rather than treating it as a strict binary classification problem. In Kalasanty 70 and PUResNet 69, each voxel is considered part of a binding pocket if it contains at least one cavity pseudoatom. The number of pseudoatoms per voxel is then clipped to the range $[0,1]$, and the model is trained with a Dice loss function as part of a binary classification task: